# Relative Preference Optimization: Enhancing LLM Alignment through Contrasting Responses across Identical and Diverse Prompts

### v1

@@ Abstract
In the field of large language models (LLMs), aligning models with the diverse preferences of users is a critical challenge. Direct Preference Optimization (DPO) has played a key role in this area. It works by using pairs of preferences derived from the same prompts, and it functions without needing an additional reward model. However, DPO does not fully reflect the complex nature of human learning, which often involves understanding contrasting responses to not only identical but also similar questions. To overcome this shortfall, we propose Relative Preference Optimization (RPO). RPO is designed to discern between more and less preferred responses derived from both identical and related prompts. It introduces a contrastive weighting mechanism, enabling the tuning of LLMs using a broader range of preference data, including both paired and unpaired sets. This approach expands the learning capabilities of the model, allowing it to leverage insights from a more varied set of prompts. Through empirical tests, including dialogue and summarization tasks, and evaluations using the AlpacaEval2.0 leaderboard, RPO has demonstrated a superior ability to align LLMs with user preferences and to improve their adaptability during the training process. The PyTorch code necessary to reproduce the results presented in the paper will be made available on GitHub [^1] https://github.com/yinyueqin/relative-preference-optimization for public access.
@@ 1 Introduction
Large language models (LLMs) like ChatGPT (OpenAI, 2023) and LLaMa (Touvron et al., 2023) have initiated a transformative era in artificial intelligence. These models exhibit remarkable abilities in fields such as natural language processing, logical reasoning, and programming (Pan et al., 2023; Tian et al., 2023). Trained on extensive, unsupervised datasets, they adeptly handle complex tasks and excel in zero-shot and few-shot learning scenarios. However, the diversity inherent in these datasets, while enriching content variety, presents alignment challenges. The vast array of human thought and expression within these datasets, embodying different goals, priorities, and values, can lead to outputs that do not consistently align with specific human preferences or desired outcomes, particularly in precision-demanding and nuanced contexts (Srivastava et al., 2022; Thoppilan et al., 2022; Bubeck et al., 2023).
To address alignment challenges, Supervised Fine-Tuning (SFT) (Wei et al., 2022; Wang et al., 2023; Yu et al., 2024) is commonly used, tailoring LLMs to tasks with labeled data. While effective, SFT may not fully capture the breadth of preference learning, essential for generating outputs aligned with human expectations beyond technical accuracy, encompassing ethical and contextual considerations. Expanding on SFT, Reinforcement Learning from Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022; Ramamurthy et al., 2023) critically aligns LLMs with human preferences. Despite its effectiveness, RLHF is resource-intensive, relying on extensive human-annotated data to train a reward model and sophisticated RL algorithms, leading to a costly and labor-intensive process. The distribution gap between model-generated outputs and training data poses a major hurdle, requiring continuous updates and re-annotations for sustained effectiveness, complicating and increasing costs associated with achieving human-aligned LLMs (Yuan et al., 2023).
@@
To address the intricacies and challenges inherent in RLHF, the Direct Preference Optimization (DPO) (Rafailov et al., 2023) approach directly fine-tunes the language model’s policy to more closely align with human preferences, thereby obviating the need for a separate and often complex reward model. Central to DPO is the utilization of pairwise preferred and rejected responses for each prompt, which form the foundation for developing a method to optimize preferences effectively. However, training a model to learn from individual preferences for each example may not fully encapsulate the complexity of human learning. Human cognition often involves interpreting divergent responses, not only to identical questions but also to those that are similar, highlighting the multifaceted nature of comprehension and preference formation (Dahlin et al., 2018). Moreover, obtaining pairwise preference data can pose challenges and incur substantial costs, particularly in sensitive domains such as healthcare and personal services, where careful attention to ethical considerations is paramount.
@@
Our inspiration draws from the human learning process, where valuable insights often arise from the juxtaposition of successful examples and relevant failures (Dahlin et al., 2018). To emulate this, we introduce Relative Preference Optimization (RPO). This approach involves analyzing prompt similarities at the semantic level within each mini-batch, allowing us to classify pairs as either highly related or unrelated. We construct a contrast matrix that instructs the model to distinguish between preferred and rejected responses, applicable to both identical and semantically related prompts. We have developed three weighting strategies to recalibrate the comparison of each contrastive pair. Our findings reveal that reweighting based on prompt similarities significantly enriches model alignment with human preferences, offering a more nuanced understanding. Furthermore, RPO inherently excels in handling non-pairwise preference data by considering semantically related contrastive pairs.
@@
As illustrated in Figure 1, we are interested in the question “Explain the concept of photosynthesis.” DPO applies penalties for incorrect responses and rewards for precise responses generated for the specific prompt. Conversely, our method RPO emphasizes the semantic connections between various prompts. For instance, the prompt “Describe the importance of sunlight in plant growth” is conceptually similar, and its responses might intersect with those of the initial question. Under RPO, a dismissed answer for the second prompt could also be seen as a misstep for the first prompt. Thus, RPO penalizes both $y_{l,1}$ and $y_{l,2}$ while approving $y_{w,1}$. It is crucial to note that not all prompts are semantically related enough to form effective contrastive pairs. RPO incorporates a reweighting mechanism, whereby unrelated prompts are given less emphasis during training. RPO expands the learning horizon of the model, empowering it to leverage insights from a broader range of prompts, mirroring the human learning process more closely.
@@
We empirically evaluate RPO on the state-of-the-art (SoTA) LLMs, such as LLama2-7/13B and Mistral-7B. We compare RPO with SoTA human preference alignment methods, and RPO significantly outperforms the baselines on typical dialogue and summarization tasks and assessments on the AlpacaEval2.0 leaderboard. The core contributions of RPO are summarized as follows:
@@
  • Innovative preference contrast learning strategy: RPO enriches the landscape of preference optimization with novel contrastive learning techniques.
@@ 2 Related Work
The field of large language models (LLMs) has made notable strides in aligning these models with human preferences, driven by innovative fine-tuning methodologies. In this exploration, we delve into key methodologies, including SFT, RLHF, DPO, and their respective variations.
2.1 Supervised Fine-Tuning (SFT)
SFT (Sanh et al., 2021; Chung et al., 2022) serves as a foundational phase in the training of LLMs. It entails customizing a pre-trained model for specific tasks with a smaller set of labeled data. The goal here is to adjust the model’s outputs to mirror the high-quality responses in the training set. The main objective of SFT revolves around reducing the negative log-likelihood loss:
$L_{\text{SFT}}(\theta)=-\mathbb{E}_{(x,y)\sim D}[\log p_{\theta}(y|x)],$ (1)
where $p_{\theta}(y|x)$ denotes the predicted probability by the model for a response $y$ to a prompt $x$, and $D$ the labeled data.
@@ 2.2 Reinforcement Learning Fine-Tuning (RLHF)
$\begin{split}p(y_{w}\succ y_{l}|x)&=\frac{\exp(r^{*}(y_{w}|x))}{\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x))}\\ &=\sigma(r^{*}(y_{w}|x)-r^{*}(y_{l}|x)),\end{split}$ (2)
@@ 2.5 Kahneman-Tversky Optimization (KTO)
Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2023) diverges from the preference likelihood maximization used in DPO. Instead, KTO focuses on maximizing the utility of model outputs, informed by the human value function derived from Kahneman-Tversky’s prospect theory. This adaptation to language models allows KTO to operate without the necessity of preference pairs, thereby streamlining data requirements. The KTO loss function is formalized as:
@@
KTO operates in unpaired data scenarios by independently processing chosen and rejected samples. As a concurrent work, our approach employs alternative mechanisms to ensure its applicability in contexts involving unpaired data.
@@ 3 Relative Preference Optimization
The traditional DPO framework aligns language model outputs with human preferences using pairwise data, where each pair is composed of a preferred (win) and rejected (lose) sample for the same prompt. However, this approach is limited to situations where such pairwise preference data is accessible, failing to exploit the valuable comparative insights that could be derived from contrasting diverse samples across a range of prompts.
In response, our RPO framework encompasses a wider array of preference data, including non-paired samples. This development not only improves the use of existing preference data but also facilitates model training in complex scenarios where pair-wise data is not readily obtainable. More specifically, RPO integrates preference pairs derived from prompts that are semantically related but not identical, as shown in Figure 1. Through dynamic calculation of relative reward weights based on prompt similarities, our method enhances the model’s ability to learn from a wider array of human feedback, leading to better generalization and more robust preference alignment.
@@ 3.1 Contrast Matrix Construction
Paired Data Scenario. In situations where each win response is associated with a corresponding lose response from the same prompt, the contrast matrix is an $M\times M$ square matrix, where $M$ represents the total count of unique prompts from a specific batch within the dataset. Each element $c_{ij}$ within this matrix represents the contrastive score between the win response of the $i^{th}$ prompt and the lose response of the $j^{th}$ prompt. For diagonal elements where $i=j$, the score reflects the direct comparison within the same prompt, while off-diagonal elements represent the relative reward differences across distinct prompts. In this context, DPO is limited to using only the diagonal terms of the contrast matrix, while RPO takes into account all pairings within the matrix, encompassing a broader range of preference comparisons.
@@
For each win response $y_{w,i}$ and lose response $y_{l,j}$, the contrastive score $s_{ij}$ is computed as the difference in rewards associated with each response, as defined by the following equations:
$\begin{split}s_{ij}&=r_{w,i}-r_{l,j},\\ r_{w,i}&=\beta\log\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}+\beta\log Z(x_{i}),\\ r_{l,j}&=\beta\log\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}+\beta\log Z(x_{j}),\end{split}$ (10)
where $\beta$ serves as a scaling factor. In this formula, the calculation of each reward follows the DPO approach (Rafailov et al., 2023). $r_{w,i}$ denotes the reward associated with the $i^{th}$ win response, and $r_{l,j}$ signifies the reward for the $j^{th}$ lose response. By calculating the contrastive scores across the matrix, RPO enables a comprehensive evaluation of the relative preference among all potential contrastive pairs.
@@ 3.2 Weighting Strategies
Figure 3: Three weighting strategies in RPO: (a) Uniform Weighting assigns equal weight to all elements within the contrast matrix, (b) Diagonal Emphasis Weighting Strategy allocates increased weight to diagonal elements while equalizing the rest, and (c) Embedding Distance Reweighting adjusts weights according to embedding distance between prompts.
After forming the contrast matrix for each batch in RPO, we deploy diverse strategies to assign differentiated weights to each comparison pair. These weights crucially determine the relative influence of different comparison pairs in the final loss computation. We propose a strategy that reweights the contrast matrix based on the distance between prompt feature embeddings for preference learning. For distinct data configurations, we have introduced two additional, simpler weighting strategies alongside our primary Embedding Distance Reweighting Strategy, as shown in Figure 3.
@@
The incorporation of prompt similarity plays a pivotal role in contrastive analysis, applicable to both paired and unpaired datasets. This technique involves calculating the cosine distance $d=\cos(f(x_{w}),f(x_{l}))$ between the embeddings of win ($x_{w}$) and lose ($x_{l}$) prompts, effectively assessing their thematic relatedness. This similarity directly influences the unnormalized weight $\tilde{\omega}$ assigned to each pair of responses:
@@
Here, $f$ represents the model for extracting sentence embeddings, such as Sentence-T5-Large (Raffel et al., 2020), and $\tau$ acts as a temperature parameter that moderates the impact of prompt similarity on the weight. Specifically, a lower $\tau$ results in a greater variation in weights for different levels of similarity, emphasizing the contrastive scores between closely related prompts. In contrast, a higher $\tau$ leads to a more uniform distribution of weights, diminishing the disparity in influence between prompts of varying thematic similarity. The adjusted contrastive score $s_{ij}$ for a win and lose response pair is then calculated as:
$s_{ij}=\omega_{ij}\times(r_{w,i}-r_{l,j}),\leavevmode\nobreak\ \leavevmode\nobreak\ \omega_{ij}=\frac{\tilde{\omega}_{ij}}{\sum_{j^{\prime}=1}^{N}\tilde{\omega}_{ij^{\prime}}}$ (12)
where $N$ represents the mini-batch size, and each $\omega_{ij}$ is normalized within its respective mini-batch so that $\sum_{j=1}^{N}\omega_{ij}=1$. In this configuration, the sum of the weights in each row of the matrix equals one. It is noteworthy that normalizing across the columns represents another valid approach. Moreover, more intricate reweighing methodologies, inspired by concepts such as optimal transport (Peyré et al., 2017) or conditional transport (Zheng and Zhou, 2021), might offer additional refinements when tuning LLMs (Tanwisuth et al., 2023). These possibilities warrant further exploration in future research.
Following KTO (Ethayarajh et al., 2023), we assume that $Z(x)$ in Equation 10 is the same for all inputs, and we obtain
@@
This strategy ensures that contrast scores derived from thematically similar prompts are accentuated, enhancing context-sensitive preference learning. In RPO, while the Embedding Distance Reweighting Strategy serves as our primary method for analyzing the contrast matrix, we have also explored two additional strategies for specific data configurations, offering alternative ways to weight the contrast matrix.
@@
where $r_{w,i}-r_{l,j}$ follows the same formulation as described in Equation 13.
@@
where $s_{ij}$ represents the adjusted contrastive scores calculated using the three weighting strategies mentioned before. This loss function directs the model to amplify the reward for winning responses and diminish it for losing ones among both identical and semantically related prompts. The learning intensity is dynamically modulated by both the prompt-aware reweighting factor $w$ and the scaling factor $\beta$. RPO broadens the learning scope of the model, enabling it to utilize insights from a wider array of prompts, facilitating a more nuanced alignment with human preferences.

### v2

@@ Abstract
In the field of large language models (LLMs), aligning models with the diverse preferences of users is a critical challenge. Direct Preference Optimization (DPO) has played a key role in this area. It works by using pairs of preferences derived from the same prompts, and it functions without needing an additional reward model. However, DPO does not fully reflect the complex nature of human learning, which often involves understanding contrasting responses to not only identical but also similar questions. To overcome this shortfall, we propose Relative Preference Optimization (RPO). RPO is designed to discern between more and less preferred responses derived from both identical and related prompts. It introduces a contrastive weighting mechanism, enabling the tuning of LLMs using a broader range of preference data, including both paired and unpaired sets. This approach expands the learning capabilities of the model, allowing it to leverage insights from a more varied set of prompts. Through empirical tests, including dialogue and summarization tasks, and evaluations using the AlpacaEval2.0 leaderboard, RPO has demonstrated a superior ability to align LLMs with user preferences and to improve their adaptability during the training process. Our code can be viewed at https://github.com/yinyueqin/relative-preference-optimization.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2023) and LLaMA (Touvron et al., 2023) have revolutionized AI, demonstrating remarkable capabilities in natural language processing, logical reasoning, and programming (Pan et al., 2023; Tian et al., 2023). Their proficiency in zero-shot and few-shot learning is attributed to training on extensive, unsupervised datasets. However, the diverse nature of these datasets can result in alignment challenges, leading to outputs that may not consistently align with specific human values, particularly in nuanced contexts (Agrawal et al., 2023; Shi et al., 2023; Liang et al., 2021; Sheng et al., 2019; Kadavath et al., 2022; Srivastava et al., 2022; Thoppilan et al., 2022; Bubeck et al., 2023). The Direct Preference Optimization (DPO) method fine-tunes the language model’s policy to align more closely with human preferences, thereby eliminating the need for a separate reward model, a staple in traditional Reinforcement Learning from Human Feedback (RLHF) (Schulman et al., 2017). Central to DPO is the utilization of pairwise preferences, with preferred and dispreferred responses identified for each prompt. This forms the foundation for effectively optimizing preferences. However, training a model to learn from an individual preference pair for each example may not fully capture the complexity of human learning. Human cognition often involves interpreting divergent responses, not only to identical questions but also to similar ones, highlighting the multifaceted nature of comprehension and preference formation (Dahlin et al., 2018). Moreover, obtaining pairwise preference data can pose challenges and incur substantial costs, especially in sensitive domains such as healthcare and personal services, where careful attention to ethical considerations is essential (Murtaza et al., 2023).
@@
Our inspiration draws from the human learning process, where valuable insights often arise from the comparison of successful examples and relevant failures (Dahlin et al., 2018). To emulate this, we introduce Relative Preference Optimization (RPO). This approach involves analyzing prompt similarities at the semantic level within each mini-batch, allowing us to classify pairs as either highly related or unrelated. We construct a contrast matrix that instructs the model to distinguish between preferred and dispreferred responses, applicable to both identical and semantically related prompts. We have developed three weighting strategies to recalibrate the comparison of each contrastive pair. Our findings reveal that reweighting based on prompt similarities significantly enriches model alignment with human preferences, offering a more nuanced understanding. Furthermore, RPO inherently excels in handling non-pairwise preference data by considering semantically related contrastive pairs.
@@
As illustrated in Figure 1, we are interested in the question “Explain the concept of photosynthesis.” DPO applies penalties for incorrect responses and rewards for precise responses generated for the specific prompt. Conversely, our method RPO emphasizes the semantic connections between various prompts. For instance, the prompt “Describe the importance of sunlight in plant growth” is conceptually similar, and its responses might intersect with those of the initial question. Under RPO, if an answer is less preferred for the second prompt, it is also treated as less suitable for the first prompt. Thus, RPO penalizes both $y_{l,1}$ and $y_{l,2}$ while approving $y_{w,1}$. It is crucial to note that not all prompts are semantically related enough to form effective contrastive pairs. RPO incorporates a reweighting mechanism, whereby unrelated prompts are given less emphasis during training. RPO expands the learning horizon of the model, empowering it to leverage insights from a broader range of prompts, mirroring the human learning process more closely.
@@
We empirically evaluate RPO on three LLMs, LLaMA2-7/13B and Mistral-7B. We compare RPO with SoTA human preference alignment methods, and RPO significantly outperforms the baselines on typical dialogue and summarization tasks and evaluations on the AlpacaEval2.0 leaderboard. The core contributions of RPO are summarized as follows:
@@
  • Innovative contrastive preference learning strategy: RPO enriches the landscape of preference optimization with novel contrastive learning techniques.
@@ 2 Related Work
The field of large language models (LLMs) has made notable strides in aligning these models with human preferences (Chung et al., 2022; Schulman et al., 2017; Zhao et al., 2023; Rafailov et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Cheng et al., 2023; Pal et al., 2024), driven by innovative fine-tuning methodologies. In this exploration, we discuss several works closely related to our research.
@@ 2.1 Reinforcement Learning Fine-Tuning (RLHF)
$\begin{split}p(y_{w}\succ y_{l}|x)=\frac{\exp(r^{*}(y_{w}|x))}{\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x))}=\sigma(r^{*}(y_{w}|x)-r^{*}(y_{l}|x)),\end{split}$ (1)
@@ 2.4 Kahneman-Tversky Optimization (KTO)
Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) diverges from the preference likelihood maximization used in DPO. Instead, KTO focuses on maximizing the utility of model outputs, informed by the human value function derived from Kahneman-Tversky’s prospect theory. This adaptation to language models allows KTO to operate without the necessity of preference pairs, thereby streamlining data requirements. The KTO loss function is formalized as:
@@
The function $w(y)$ represents the weighting factor applied to the KTO loss function. By default, $w(y)=1$. KTO operates in unpaired data scenarios by independently processing chosen and rejected samples. As a concurrent work, our approach employs alternative mechanisms to ensure its applicability in contexts involving unpaired data.
@@ 3 Relative Preference Optimization
The traditional DPO framework aligns language model outputs with human preferences using pairwise data, where each pair is composed of a preferred (win) and dispreferred (lose) sample for the same prompt. However, this approach is limited to situations where such pairwise preference data is accessible, failing to exploit the valuable comparative insights that could be derived from contrasting diverse samples across a range of prompts. In response, our RPO framework encompasses a wider array of preference data, including non-paired samples. This development not only improves the use of existing preference data but also facilitates model training in complex scenarios where pair-wise data is not readily obtainable. More specifically, RPO integrates preference pairs derived from prompts that are semantically related but not identical, as shown in Figure 1. Through dynamic calculation of relative reward weights based on prompt similarities, our method enhances the model’s ability to learn from a wider array of human feedback, resulting in better preference alignment.
@@ 3.1 Contrast Matrix Construction
Paired Data Scenario. In situations where each win response is associated with a corresponding lose response from the same prompt, the contrast matrix is an $M\times M$ square matrix, where $M$ represents the total count of unique prompts from a specific mini-batch within the dataset. Each element $c_{ij}$ within this matrix represents the contrastive score between the win response of the $i^{th}$ prompt and the lose response of the $j^{th}$ prompt. For diagonal elements where $i=j$, the score reflects the direct comparison within the same prompt, while off-diagonal elements represent the relative reward differences across distinct prompts. In this context, DPO is limited to using only the diagonal terms of the contrast matrix, while RPO takes into account all pairings within the matrix, encompassing a broader range of preference comparisons.
@@
For each win response $y_{w,i}$ and lose response $y_{l,j}$, the contrastive score $s_{ij}$ is computed as the difference in rewards associated with each response, defined by the following equations:
$s_{ij}=r_{w,i}-r_{l,j},~{}~{}\begin{cases}r_{w,i}=\beta\log\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}+\beta\log Z(x_{i}),\\ r_{l,j}=\beta\log\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}+\beta\log Z(x_{j}),\end{cases}$ (9)
where $r_{w,i}$ denotes the reward associated with the $i^{th}$ win response, and $r_{l,j}$ signifies the reward for the $j^{th}$ lose response. By calculating the contrastive scores across the matrix, RPO enables a comprehensive evaluation of the relative preference among all potential contrastive pairs.
Similar to DPO (Rafailov et al., 2023), we define $Z(x)=\sum_{y}[\pi_{ref}(y|x)\exp(\frac{1}{\beta}\cdot r(x,y))]$. The normalization term $Z(x)$ in DPO is considered a nuanced constant that could generally be omitted. However, with the introduction of cross-prompt contrast in RPO, differences between $Z(x_{i})$ and $Z(x_{j})$ become meaningful, illustrating how well the reference model responds to prompts $x_{i}$ versus $x_{j}$. This variance offers opportunities to differentiate prompts based on their responses from the model. Nevertheless, estimating $Z(x)$ accurately remains a challenge. As detailed in Appendix A, it is reasonable to assume that differences between $Z(x_{i})$ and $Z(x_{j})$ correlate with the distinctions between $x_{i}$ and $x_{j}$. The weighting strategy we discuss later allows us to reasonably disregard the differences between $Z(x_{i})$ and $Z(x_{j})$ when computing $s_{ij}$.
@@ 3.2 Weighting Strategies
After forming the contrast matrix for each mini-batch in RPO, we deploy diverse strategies to assign differentiated weights to each comparison pair. These weights crucially determine the relative influence of different comparison pairs in the final loss computation. We propose a strategy that reweights the contrast matrix based on the distance between prompt feature embeddings for preference learning. For distinct data configurations, we have introduced two additional, simpler weighting strategies alongside our primary Embedding Distance Reweighting Strategy.
@@
The incorporation of prompt similarity plays a pivotal role in contrastive analysis, applicable to both paired and unpaired datasets. This technique involves calculating the cosine distance $d=\cos(f(x_{w}),f(x_{l}))$ between the embeddings of win ($x_{w}$) and lose ($x_{l}$) prompts, effectively assessing their thematic relatedness. This similarity directly influences the weight $\tilde{\omega}$ assigned to each pair of responses:
@@
Here, $f$ represents the model for extracting sentence embeddings, such as all-MiniLM-L6-v2 (Wang et al., 2020), and $\tau$ acts as a temperature parameter that moderates the impact of prompt similarity on the weight. Specifically, a lower $\tau$ results in a greater variation in weights for different levels of similarity, emphasizing the contrastive scores between closely related prompts. In contrast, a higher $\tau$ leads to a more uniform distribution of weights, diminishing the disparity in influence between prompts of varying thematic similarity. The adjusted contrastive score $s_{ij}$ for a win and lose response pair is then calculated as:
$s_{ij}=\omega_{ij}\times(r_{w,i}-r_{l,j}),~{}~{}\omega_{ij}=\frac{\tilde{\omega}_{ij}}{\sum_{j^{\prime}=1}^{N}\tilde{\omega}_{ij^{\prime}}}$ (11)
where $N$ represents the number of lose responses in the mini-batch, and each $\omega_{ij}$ is normalized within its respective mini-batch so that $\sum_{j=1}^{N}\omega_{ij}=1$. In this configuration, the sum of the weights in each row of the matrix equals one.
We then obtain
@@
This strategy ensures that contrast scores derived from thematically similar prompts are accentuated, enhancing context-sensitive preference learning. While the Embedding Distance Reweighting Strategy serves as the primary method to construct RPO’s contrast matrix, we have explored two additional strategies for specific data configurations, offering alternative ways to weight the contrast matrix.
@@
where $r_{w,i}-r_{l,j}$ follows the same formulation as described in Eq. 12.
@@
where $s_{ij}$ represents the adjusted contrastive scores calculated using one of the three weighting strategies mentioned before. This loss function directs the model to amplify the reward for winning responses and diminish it for losing ones among both identical and semantically related prompts. The learning intensity is dynamically modulated by both the prompt-aware reweighting factor $w$ and the scaling factor $\beta$.